### v2

@@ Abstract
Coordinated online behavior, which spans from beneficial collective actions to harmful manipulation such as disinformation campaigns, has become a key focus in digital ecosystem analysis. Traditional methods often rely on monomodal approaches, focusing on single types of interactions like co-retweets or co-hashtags, or consider multiple modalities independently of each other. However, these approaches may overlook the complex dynamics inherent in multimodal coordination. This study compares different ways of operationalizing the detection of multimodal coordinated behavior. It examines the trade-off between weakly and strongly integrated multimodal models, highlighting the balance between capturing broader coordination patterns and identifying tightly coordinated behavior. By comparing monomodal and multimodal approaches, we assess the unique contributions of different data modalities and explore how varying implementations of multimodality impact detection outcomes. Our findings reveal that not all the modalities provide distinct insights, but that with a multimodal approach we can get a more comprehensive understanding of coordination dynamics. This work enhances the ability to detect and analyze coordinated online behavior, offering new perspectives for safeguarding the integrity of digital platforms.
@@ 1 Introduction
The digital age has transformed the way individuals and organizations interact online, giving rise to novel forms of coordination in social and political arenas. Coordinated behavior, a phenomenon where multiple actors engage in synergic actions in pursuit of an intent [1], has become increasingly relevant in the study of online ecosystems. Coordination in online behavior can manifest across a wide spectrum of intentions and outcomes. It may emerge organically or be deliberately orchestrated, and its perceived value often depends on the observer’s perspective. For instance, coordinated efforts are found in activism [2, 3], boycotts [4, 5], and protests [6, 7], which some may view as socially beneficial, while others may see them as disruptive or even harmful. Conversely, coordination can also be employed for goals widely recognized as problematic, such as disinformation campaigns [8, 9, 10], election interference [3, 11], or astroturfing [12]. Even in cases like coordinated botnets [13, 14, 15, 16], which are typically associated with manipulation, there may be applications aligned with legitimate objectives. Ultimately, the evaluation of coordinated behavior – whether deemed harmful or beneficial – is often subjective and context-dependent. Hence, since coordination in social media exhibits a dual nature, the scope behind coordination can vary widely, ranging from benign collective action to malicious intent aimed at manipulation or deception.
[…]
Refer to captionRefer to captionco-favorite
$w_{fav}$
@@
Understanding and detecting coordinated online behavior is critical for maintaining the integrity of digital platforms [17]. Researchers and policymakers alike seek to develop methodologies to identify patterns of coordination, disentangle genuine collective action from inauthentic manipulation, and mitigate the harmful consequences of coordinated inauthentic behavior. Various computational techniques, including network-based analysis [18, 2, 6, 19] and machine learning models [20, 21, 22] have been employed to unveil hidden networks of coordination. Users may engage in a wide range of coordinated actions, which complicates their detection. For example, users might coordinate through synchronized retweets to amplify specific content [11, 23, 24], through likes [25], or by using identical hashtags to dominate trending topics [26, 27]. More covert forms of coordination can include strategic timing of posts [28], the use of similar narratives across multiple accounts, or the deliberate targeting of specific audiences. These varied manifestations highlight the inherent multimodal nature of coordination in social media, where the mode of action plays a crucial role in shaping the dynamics and outcomes of coordinated behavior. Figure 1 shows how coordinated online behavior is inherently multimodal, as users can coordinate through different types of actions. Therefore, it is crucial to design and provide a robust methodology capable of uncovering coordinated behaviors that manifest through diverse user actions.
@@
While network-based methods have proven effective in detecting such behavior, most approaches either focus on a single type of action or treat different actions separately, resulting in the construction of multiple single-layer networks that miss the opportunity to analyze them in an integrated way. Although some studies have begun to explore multimodal interactions, they fail to fully leverage the richness and complexity that such diversity offers [29, 6]. Many of these approaches flatten the layers into a single aggregated network, thereby losing the distinct characteristics of each modality and the interplay between them [30, 31, 26, 2]. The current landscape reveals two key limitations: first, existing multimodal analyses do not truly exploit the potential of integrating different types of actions; second, there is a lack of systematic investigation into how and whether multimodality offers advantages over analyzing each type of action independently. As a result, we lack not only a clear understanding of the value of multimodality in this context, but also effective frameworks or methods to operationalize it properly. Addressing these gaps is essential for advancing our ability to detect and interpret coordinated behavior in a more comprehensive and meaningful way.
@@
For this reason, it is necessary to adopt a comprehensive approach that accounts for the multimodal nature of coordinated behavior. In this work, we propose different ways to operationalize the detection of multimodal coordinated behavior, by incorporating various types of actions and, consequently, different forms of coordination. By integrating multiple modalities – such as retweets, mentions, hashtag usage, and posting patterns – our method provides a holistic perspective on how users coordinate their efforts. This approach not only captures the diversity of coordinated actions but also enables the detection of subtle and complex behaviors that might be overlooked when focusing on a single modality. By bridging these different dimensions of coordination, our method enhances the understanding and identification of coordinated efforts in social media, contributing to more robust and reliable analysis frameworks.
@@
Our framework includes several multimodal detection methods for coordinated behavior and Figure 2 presents examples of possible ways to operationalize multimodality, along with the number of studies in the literature that have adopted each approach. For example, the method frequently used in the literature – flattening the layers into a single network – could be one solution, albeit perhaps not the most effective. Another approach, which considers different coordination networks and actions independently, as is often done, represents the weakest form of multimodality operationalization, as it overlooks the interplay between modalities. On the other end of the spectrum, the strongest approach forces the model to recognize users as coordinated only if they demonstrate coordination across all types of actions. This stricter operationalization fully embraces multimodality but at the cost of increased selectivity. As we can anticipate, there is an inherent tradeoff in leveraging multimodality. While enriching the model by integrating more modalities allows for the detection of complex and nuanced forms of coordinated behavior, increasing the intersection of multimodal conditions too much may overly constrain the analysis, potentially limiting the detection to only a few instances of coordination. Figure 2 illustrates this tradeoff, highlighting the balance between the level of multimodality and constraints loosening in coordinated behavior detection.
lowhighmultimodalityhighlowconstraints loosening monomodal independent layers union flattening multiplex community detection intersection flattening >20 papers >15 papers >4 papers zero papers zero papers
@@
In summary, our approach not only advances the detection of multimodal coordinated behavior but also provides a comprehensive exploration of how to effectively operationalize multimodality, striking a balance between complexity and detectability.
@@ 1.1 Research focus
  • RQ2 How do multimodal approaches contribute to the analysis of coordinated online behavior compared to monomodal approaches? Additionally, how do different implementations of multimodality affect the results? Here we explore the added value of multimodal approaches in detecting and analyzing coordinated online behavior compared to monomodal analyses. While monomodal approaches focus on a single type of interaction, multimodal approaches leverage multiple modalities simultaneously, potentially offering a more comprehensive view of coordination patterns. However, multimodality is a flexible concept that encompasses various ways of combining modalities. Different strategies for integrating modalities may highlight different aspects of coordinated behavior and yield varying analytical outcomes. This question thus addresses two key points: the overall benefit of multimodal analyses relative to monomodal ones, and the impact that different implementations of multimodality have on the results.
To address the research questions outlined above, we conduct our analysis on a Twitter dataset. This dataset offers a rich and appropriate context for studying coordinated online behavior due to the platform’s public and interaction-driven nature. It has been previously used in the study of coordination dynamics, although not in a multimodal setting [3, 32, 23, 33].
@@ 1.2 Contributions
  • We propose the first comparative analysis of operationalizations for detecting multimodal coordinated behavior, including an approach that fully leverages the multimodal structure of a multiplex coordination network and capturing interdependencies among different modalities through a community detection algorithm specifically designed for multiplex structures.
@@ 2 Related work
Coordination detection methods can be classified into two main categories depending on their underlying approach: network science or machine learning. Network science approaches can be categorized in single-layer or multi-layer ones. The following sections discuss the existing solutions in each category.
@@ 2.1 Machine learning-based approaches
This category of methods applies data mining and machine learning techniques to extract pattern and insights from various data types, including text [34, 35, 36], images [37], audio [22], interactions [20, 21, 7, 38], and temporal data [39, 40, 41]. However, only a few of these approaches extract multimodal features from the data or leverage multiple types of user activity. Despite the diversity of machine learning methods – and the absence of a unifying framework comparable to that of network-based approaches – they can generally be grouped into two broad categories: unsupervised methods, which cluster users based on behavioral similarities, and supervised methods, which classify users as coordinated or non-coordinated based on labeled examples.
Among unsupervised approaches, some methods model user activity as temporal point processes. For instance, Sharma et al. [20] introduced AMDN-HAGE, a generative model that combines activity traces (AMDN) with user group estimation (HAGE) to detect synchronized user groups – assuming all detected coordination to be malicious. Similarly, Zhang et al. [21] jointly learn user-group distributions by leveraging embeddings and temporal logic, clustering users through an expectation-maximization algorithm. In a different direction, Erhardt and Albassam [42], Erhardt and Pentland [43] represent user behaviors as interacting Markov chains within a discrete-time stochastic model to capture coordination dynamics.
Supervised methods, on the other hand, rely on labeled data to identify coordinated users. Zhang et al. [38] address cross-platform coordination by using an auxiliary platform with known coordinated users to train neural time series encoders, followed by a multi-layer perceptron to detect similar behavior on a target platform. Mariconti et al. [22] detect coordinated hate attacks on YouTube by extracting features from video metadata, audio, and thumbnails, and feeding them into an ensemble classifier trained on content that had previously been targeted by coordinated raids. A more lightweight approach is proposed by Francois et al. [41], who construct a small ground-truth of organic and inorganic campaigns and classify new ones based on the confidence intervals of network, temporal, and semantic indicators. ì
@@ 2.2 Network science-based single-layer approaches
Most of detection methods in literature exploit the so-called coordination network, a graph where nodes represent users and edges represent a co-action, i.e., an action of the same type performed by two users. For instance, two users retweeting the same tweet perform a co-retweet, or posting a post with the same hashtag, a co-hashtag. Generally, the weight of an edge is proportional to the number of co-actions between the two users. Communities of users can be detected in the coordination network to identify groups of users that are more densely connected to each other than to the rest of the network. These communities are interpreted as groups of users that are coordinating their actions [1]. The majority of the works in literature are monomodal approaches, as they focus on a single type of co-action, such as co-retweet [18, 3, 11, 44], co-hashtag [45, 27, 46], co-mention [47], and co-like [25]. Many others construct several coordination network, one for each co-action, to perform independent analysis and comparing the results. For instance, Weber and Neumann [48, 19] analyze co-retweet, co-URL, co-hashtag, co-mention, and co-reply, Yu [49] analyze users posting posts with similar images and videos, and Ng and Carley [30] construct a co-URL and co-tweet coordination network, i.e., respectively, users posting similar posts and posts with the same URLs.
Although, they perform analysis on different types of co-actions, they do not consider the interdependence between them. Indeed, a model based on the construction of a coordination network for each type of co-action does not capture the interplay between different types of co-actions. For instance, Dash and Mitra [50] refers to their analysis as a multimodal analysis, even if they detect the coordinated communities in each layer independently. With respect to Figure 2, this work falls within the independent layers approach.
@@ 2.3 Network science-based multi-layer approaches
As already discussed, online coordination is a complex phenomenon that can involve different types of actions, possibly simultaneously. In this context, some works have proposed to model the coordinated behavior as a multiplex coordination network [29, 6]. At a high level, a multiplex network is a type of network that consists of multiple layers, where each layer represents a different type of relationship or interaction between the same set of nodes. In the context of coordinated behavior, each layer represents a different type of co-action, and the multiplex network can capture the interplay between different types of co-actions. However, modeling the coordination as a multiplex network is not sufficient to capture the multimodality of coordinated behavior if the method used to detect coordinated communities does not fully exploit the multiplex nature of the network [31, 51, 26, 2, 30, 52].
@@
Indeed, most of the works declaring to exploit a multiplex network, in fact, flattens the multiplex network into a single-layer network before applying the community detection method – union flattening in Figure 2. This is done by summing the weights of the edges in the different layers [2], by considering the union of the edges in the different layers [30, 31, 26], or by creating a multigraph, where all the layers are merged into a single layer [52]. This flattening step is motivated by the fact that most of the community detection methods are designed to work on single-layer networks, and are not able to handle multiplex networks. However, this flattening step may lead to a loss of information, as it does not take into account the different types of co-actions.
@@
A couple of works leverages a multi-view clustering algorithm, which is able to handle multiplex networks without flattening them [29, 6]. However, the leveraged algorithm, proposed by Cruickshank [53], optimizes the modularity in each layer, not considering the inter-layer modularity. This means that the clusters are not allowed to span between views or layers, and the modularity can be reduced to just the sum of the intra-layer modularities. This approach is different from the method proposed in this work, which allows communities to span multiple layers. Finally, a work by Emeric and Victor [51] proposes a method to detect coordinated behavior in multiplex networks by applying Louvain on each layer and an iterative probabilistic voting consensus algorithm to achieve consensus clustering [51]. Also this work does not allow communities to span multiple layers and firstly optimizes each layer indipendently.
@@
Finally, the main drawback of the aforementioned works is not only the lack of a method that can fully exploit the multimodal nature of a multiplex coordination network, but also the lack of a comprehensive definition and evaluation of what multimodal coordinated behavior is and how its detection can be operationalized. To address this gap, we present the first comparative analysis of different ways to operationalize the detection of multimodal coordinated behavior, including an approach that fully leverages the multimodal structure of a multiplex coordination network. We also investigate the contribution of the multimodality compared to the monomodal approaches. Indeed, it is still an open question whether different modalities of co-actions provide distinct information for the analysis of coordinated behavior, or they yield equivalent results in terms of coordinated community discovered.

### v3

@@ Abstract
Coordinated online behavior, which spans from beneficial collective actions to harmful manipulation such as disinformation campaigns, has become a key focus in digital ecosystem analysis. Traditional methods often rely on monomodal approaches, focusing on single types of interactions like co-retweets or co-hashtags, or consider multiple modalities independently of each other. However, these approaches may overlook the complex dynamics inherent in multimodal coordination. This study compares different ways of operationalizing multimodal coordinated behavior, examining the trade-off between weakly and strongly integrated models and their ability to capture broad versus tightly aligned coordination patterns. By contrasting monomodal, flattened, and multimodal methods, we evaluate the distinct contributions of each modality and the impact of different integration strategies. Our findings show that while not all modalities provide unique insights, multimodal analysis consistently offers a more informative representation of coordinated behavior, preserving structures that monomodal and flattened approaches often lose. This work enhances the ability to detect and analyze coordinated online behavior, offering new perspectives for safeguarding the integrity of digital platforms.
@@
{highlights}
We introduce an operationalization to detect multimodal coordinated behavior, which fully exploits the potential of the interdependence among different modalities (e.g., co-retweets, co-hashtags, co-URLs), by exploiting a multiplex coordination network.
We analyze the contributions of different monomodal approaches in detecting coordinated online behavior, revealing whether various data modalities provide complementary or overlapping insights.
We compare different levels of multimodal integration, highlighting the trade-offs between broadly capturing coordination patterns and strictly identifying tightly coordinated users, providing guidance for robust multimodal analyses.
footnotetext: 0000-0002-5556-3746 (Lorenzo Mannocci), 0000-0003-0170-2445 (Stefano Cresci), 0000-0002-3437-9018 (Matteo Magnani), 0000-0001-8541-0284 (Anna Monreale), 0000-0001-8228-7807 (Maurizio Tesconi)
@@ 1 Introduction
The digital age has transformed the way individuals and organizations interact online, giving rise to novel forms of coordination in social and political arenas. Coordinated behavior, a phenomenon where multiple actors engage in synergic actions in pursuit of an intent [1], has become increasingly relevant in the study of online ecosystems.
Coordination in online behavior can manifest across a wide spectrum of intentions and outcomes. It may emerge organically or be deliberately orchestrated, and its perceived value often depends on the observer’s perspective. For instance, coordinated efforts are found in activism [2, 3], boycott [4], and protests [5, 6], which some may view as socially beneficial, while others may see them as disruptive or even harmful. Conversely, coordination can also be employed for goals widely recognized as problematic, such as disinformation campaigns [7], election interference [3], or astroturfing [8]. Even in cases like coordinated botnets [9], which are typically associated with manipulation, there may be applications aligned with legitimate objectives. Ultimately, the evaluation of coordinated behavior – whether deemed harmful or beneficial – is often subjective and context-dependent. Hence, since coordination in social media exhibits a dual nature, the scope behind coordination can vary widely, ranging from benign collective action to malicious intent aimed at manipulation or deception.
@@
Understanding and detecting coordinated online behavior is critical for maintaining the integrity of digital platforms. Researchers and policymakers alike seek to develop methodologies to identify patterns of coordination, disentangle genuine collective action from inauthentic manipulation, and mitigate the harmful consequences of coordinated inauthentic behavior.
@@
Detection approaches. Various computational techniques, including network-based analysis [10, 11] and machine learning models [12, 13] have been employed to unveil hidden networks of coordination. Users may engage in a wide range of coordinated actions, which complicates their detection. For example, users might coordinate through synchronized retweets to amplify specific content [14], through likes, or by using identical hashtags to dominate trending topics [15, 16]. More covert forms of coordination can include strategic timing of posts [17], the use of similar narratives across multiple accounts, or the deliberate targeting of specific audiences. These varied manifestations highlight the inherent multimodal nature of coordination in social media, where the modality of the actions plays a crucial role in shaping the dynamics and outcomes of coordinated behavior. Figure 1 shows how coordinated online behavior is inherently multimodal, as users can coordinate through different types of actions. Therefore, it is crucial to design and provide a robust methodology capable of uncovering coordinated behaviors that manifest through diverse user actions.
@@
Network-based approaches and multimodality. While network-based methods have proven effective in detecting such behavior, most approaches either focus on a single type of action or treat different actions separately, resulting in the construction of multiple single-layer networks that miss the opportunity to analyze them in an integrated way. Although some studies have begun to explore multimodal interactions, they fail to fully leverage the richness and complexity that such diversity offers [18, 5]. Many of these approaches flatten the layers into a single aggregated network, thereby losing the distinct characteristics of each modality and the interplay between them [19, 15]. The current landscape reveals two key limitations: first, existing multimodal analyses do not truly exploit the potential of integrating different types of actions; second, there is a lack of systematic investigation into how and whether multimodality offers advantages over analyzing each type of action independently. As a result, we lack not only a clear understanding of the value of multimodality in this context but also effective frameworks or methods to operationalize it properly. Addressing these gaps is essential for advancing our ability to detect and interpret coordinated behavior in a more comprehensive and meaningful way.
@@
Our work. For this reason, it is important to assess the potential benefits of adopting a comprehensive approach that accounts for the multimodal nature of coordinated behavior. In this work, we propose different ways to operationalize the detection of multimodal coordinated behavior by incorporating various types of actions and, consequently, different forms of coordination. By integrating multiple modalities – such as retweets, mentions, hashtag usage, and posting patterns – these operationalizations provide a holistic perspective on how users coordinate their efforts. The objective is not only to capture the diversity of coordinated actions but also to detect subtle and complex behaviors that might be overlooked when focusing on a single modality. By bridging these different dimensions of coordination, we enable an enhanced understanding and identification of coordinated efforts in social media, contributing to more robust and reliable analysis frameworks.
@@
Multimodality trade-off. Our framework includes several multimodal detection approaches for coordinated behavior, and Figure 2 presents examples of possible ways to operationalize multimodality, along with the number of studies in the literature that have adopted each approach. For example, the approach frequently used in the literature – flattening the layers into a single network – could be one solution, albeit perhaps not the most effective. Another approach, which considers different coordination networks and actions independently, as is often done, represents the weakest form of multimodality operationalization, as it overlooks the interplay between modalities. On the other end of the spectrum, the strongest approach forces the model to recognize users as coordinated only if they demonstrate coordination across all types of actions. This stricter operationalization fully embraces multimodality but at the cost of increased selectivity. As we can anticipate, there is an inherent trade-off in leveraging multimodality. While enriching the model by integrating more modalities allows for the detection of complex and nuanced forms of coordinated behavior, increasing the intersection of multimodal conditions too much may overly constrain the analysis, potentially limiting the detection to only a few instances of coordination. Figure 2 illustrates this trade-off, highlighting the balance between the level of multimodality and constraints loosening in coordinated behavior detection.
@@
Main findings. Experiments conducted on two Twitter/X datasets demonstrate the potential benefit of using our multimodal framework in capturing coordinated behavior across diverse types of user actions. By contrasting monomodal and multimodal analyses, we show that relying on a single co-action can obscure important patterns, whereas integrating multiple actions reveals a richer and more complete picture of coordination. Identifying in advance which co-action is most informative remains challenging, further motivating the need for multimodality to avoid overlooking meaningful signals. Our results also highlight that flattening the network may lead to a substantial loss of structural information, reducing the ability to detect coordinated communities. In contrast, the multimodal approach preserves nearly all communities identified by monomodal methods while also uncovering additional structures that would otherwise remain hidden.
In summary, this study not only expands the toolbox of methods for the detection of multimodal coordinated behavior, but also provides a comprehensive exploration of how to effectively operationalize multimodality, striking a balance between complexity and detectability.
@@ 1.1 Research focus
  • RQ2 How do multimodal approaches contribute to the analysis of coordinated online behavior compared to monomodal approaches? Additionally, how do different implementations of multimodality affect the results? Here, we explore the added value of multimodal approaches in detecting and analyzing coordinated online behavior compared to monomodal analyses. While monomodal approaches focus on a single type of interaction, multimodal approaches leverage multiple modalities simultaneously, potentially offering a more comprehensive view of coordination patterns. However, multimodality is a flexible concept that encompasses various ways of combining modalities. Different strategies for integrating modalities may highlight different aspects of coordinated behavior and yield varying analytical outcomes. This question thus addresses two key points: the overall benefit of multimodal analyses relative to monomodal ones, and the impact that different implementations of multimodality have on the results.
To address the research questions outlined above, we conduct our analysis on two Twitter datasets. One dataset offers a rich and appropriate context for studying coordinated online behavior due to the platform’s public and interaction-driven nature. It has been previously used in the study of coordination dynamics, although not in a multimodal setting [3, 20, 14]. Conversely, the second dataset includes a ground-truth labeling, enabling a validation of the ability of multimodal approaches not only to identify more structures, but also that these structures capture actual coordination.
While coordinated behavior can be evaluated through various downstream tasks, the focus of this work is methodological: we aim to compare network science–based multimodal formulations and assess how each data modality contributes to detecting patterns associated with coordinated behavior. The coordinated communities identified through our methods could be used in downstream analyses – such as studying propaganda dissemination [20], political leaning [3], or bot likelihood [20]. However, these analyses lie beyond the scope of this work. Our aim is to systematically investigate the added value of multimodality within network-based frameworks for coordinated behavior detection.
@@ 1.2 Contributions
  • We perform the first comparative analysis of operationalizations for detecting multimodal coordinated behavior, including an approach that fully leverages the multimodal structure of a multiplex coordination network and captures interdependencies among different modalities through a community detection algorithm specifically designed for multiplex structures.
@@ 2 Related work
Coordination detection methods can be classified into two main categories depending on their underlying approach: network science or machine learning. Network science approaches can be categorized into single-layer or multi-layer ones. The following sections discuss the existing solutions in each category.
@@ 2.1 Machine learning-based approaches
This category of methods applies data mining and machine learning techniques to extract patterns and insights from various data types, including text [21], images [22], audio [23], and interactions [12, 13]. However, only a few of these approaches extract multimodal features from the data or leverage multiple types of user activity. Despite the diversity of machine learning methods – and the absence of a unifying framework comparable to that of network-based approaches – they can generally be grouped into two broad categories: unsupervised methods, which cluster users based on behavioral similarities, and supervised methods, which classify users as coordinated or non-coordinated based on labeled examples.
Among unsupervised approaches, some methods model user activity as temporal point processes. For instance, Zhang et al. [12] introduced AMDN-HAGE, a generative model that combines activity traces (AMDN) with user group estimation (HAGE) to detect synchronized user groups – assuming all detected coordination to be malicious. Similarly, Zhang et al. [13] jointly learn user-group distributions by leveraging embeddings and temporal logic, clustering users through an expectation-maximization algorithm. In a different direction, Erhardt and Albassam [24], Erhardt and Pentland [25] represent user behaviors as interacting Markov chains within a discrete-time stochastic model to capture coordination dynamics. While Smith et al. [26] use Bayesian inference to identify groups of accounts that share similar account-level characteristics and target similar narratives.
Supervised methods, on the other hand, rely on labeled data to identify coordinated users. Zhang et al. [27] address cross-platform coordination by using an auxiliary platform with known coordinated users to train neural time series encoders, followed by a multi-layer perceptron to detect similar behavior on a target platform. Mariconti et al. [23] detect coordinated hate attacks on YouTube by extracting features from video metadata, audio, and thumbnails, and feeding them into an ensemble classifier trained on content that had previously been targeted by coordinated raids.
Ground-truth information is rarely available in coordinated behavior detection tasks, which limits the applicability of supervised learning approaches. Consequently, recent research has increasingly explored unsupervised machine learning methods – often incorporating multimodal information – to model user activity and detect coordinated patterns [27, 13, 12]. These methods typically rely on embedding the temporal activity of users into latent spaces that capture behavioral similarities. However, such approaches suffer from a lack of interpretability, making it difficult to understand or explain the mechanisms underlying detected coordination. This is a limitation, as in many cases the goal extends beyond merely identifying coordinated users to include the characterization of the behaviors and interactions that define their coordination [1].
@@ 2.2 Network science-based single-layer approaches
Most of the detection methods in the literature exploit the so-called coordination network, a graph where nodes represent users and edges represent a co-action, i.e., an action of the same type performed by two users. For instance, two users retweeting the same tweet perform a co-retweet, or posting a post with the same hashtag, a co-hashtag. Generally, the weight of an edge is proportional to the number of co-actions between the two users. Communities of users can be detected in the coordination network to identify groups of users that are more densely connected to each other than to the rest of the network. These communities are interpreted as groups of users that are coordinating their actions [1]. The majority of the works in literature are monomodal approaches, as they focus on a single type of co-action, such as co-retweet [10, 3], co-hashtag [28], and co-URL [29]. Many others construct several coordination networks, one for each co-action, to perform independent analysis and compare the results [30]. For instance, Weber and Neumann [11] analyze co-retweet, co-URL, co-hashtag, co-mention, and co-reply, and Ng and Carley [31] construct a co-URL and co-tweet coordination network, i.e., respectively, users posting similar posts and posts with the same URLs.
Although they perform analysis on different types of co-actions, they do not consider the interdependence between them. Indeed, a model based on the construction of a coordination network for each type of co-action does not capture the interplay between different types of co-actions. For instance, Dash and Mitra [32] refer to their analysis as a multimodal analysis, even if they detect the coordinated communities in each layer independently. With respect to Figure 2, this work falls within the independent layers approach.
@@ 2.3 Network science-based multi-layer approaches
As already discussed, online coordination is a complex phenomenon that can involve different types of actions, possibly simultaneously. In this context, some works have proposed to model the coordinated behavior as a multiplex coordination network [18, 5]. At a high level, a multiplex network is a type of network that consists of multiple layers, where each layer represents a different type of relationship or interaction between the same set of nodes. In the context of coordinated behavior, each layer represents a different type of co-action, and the multiplex network can capture the interplay between different types of co-actions. However, modeling the coordination as a multiplex network is not sufficient to capture the multimodality of coordinated behavior if the method used to detect coordinated communities does not fully exploit the multiplex nature of the network [19, 33, 2, 31, 34].
@@
Indeed, most of the works declaring to exploit a multiplex network, in fact, flatten the multiplex network into a single-layer network before applying the community detection method – union flattening in Figure 2. This is done by summing the weights of the edges in the different layers [2], by considering the union of the edges in the different layers [31, 19, 15], or by creating a multigraph, where all the layers are merged into a single layer [34]. This flattening step is motivated by the fact that most of the community detection methods are designed to work on single-layer networks and are not able to handle multiplex networks. However, this flattening step may lead to a loss of information, as it does not take into account the different types of co-actions.
@@
A couple of works leverage a multi-view clustering algorithm, which can handle multiplex networks without flattening them [18, 5]. However, the leveraged algorithm, proposed by Cruickshank [35], optimizes the modularity in each layer, not considering the inter-layer modularity. This means that the clusters are not allowed to span between views or layers, and the modularity can be reduced to just the sum of the intra-layer modularities. This approach is different from the method proposed in this work, which allows communities to span multiple layers. Finally, a work by Emeric and Victor [33] proposes a method to detect coordinated behavior in multiplex networks by applying Louvain on each layer and an iterative probabilistic voting consensus algorithm to achieve consensus clustering [33]. Also, this work does not allow communities to span multiple layers and first optimizes each layer independently.
@@
Finally, the main limitation of the aforementioned works is not only the lack of a method that can fully exploit the multimodal nature of a multiplex coordination network, but also the lack of a comprehensive definition and evaluation of what multimodal coordinated behavior is and how its detection can be operationalized. To address this gap, we present the first comparative analysis of different ways to operationalize the detection of multimodal coordinated behavior, including an approach that fully leverages the multimodal structure of a multiplex coordination network. We also investigate the contribution of the multimodality compared to the monomodal approaches. Indeed, it is still an open question whether different modalities of co-actions provide distinct information for the analysis of coordinated behavior, or whether they yield equivalent results in terms of coordinated community discovered.